A company has five applicants for two positions: two women and three men. Suppose that the five applicants are equally qualified and that no preference is given for choosing either gender. Let X equal to the number of women chosen to fill the two positions.
a) Write down the appropriate probability distribution with its parameters?
b) What are the mean and variance of this distribution?
The mean= 0.8
The standard deviation= 0.6
c) What is the probability that a company will hire exactly one woman?

Answer by stanbon(75887)

A company has five applicants for two positions: two women and three men. Suppose that the five applicants are equally qualified and that no preference is given for choosing either gender. Let X equal to the number of women chosen to fill thbe two positions.
a) Write down the appropriate probability distribution with its parameters?
binomial probability distribution
----------------------------------
b) What are the mean and variance of this distribution?
The mean= n*p = 2*(2/5) = 4/5 = 0.8
-----------------------
The standard deviation= sqrt(n*p*q) = sqrt(.8(3/5))) = 0.6928
---------
c) What is the probability that a company will hire exactly one woman?
P(x = 1) = binompdf(2,2/5,1) = 0.48
------------
Cheers,
Stan H.
------------